OpenStudy (janu16):

Given the function f(x)=4(x+3)-5, solve for the inverse function when x=3.

OpenStudy (janu16):

plz help

OpenStudy (janu16):

@Mehek14

OpenStudy (janu16):

@Michele_Laino

OpenStudy (anonymous):

To find the inverse of a function, you always want to change x and y, then solve for y again. First, change f(x) to y to avoid any confusion. \[y=4(x+3)-5\] Now, swap x and y. \[x=4(y+3)-5\] All you have to do now is solve for y. Once you've done that, substitute 3 in for x, and you'll be all set!

OpenStudy (janu16):

Wait so how do you substitute x for 3?

OpenStudy (janu16):

@adrianna45

OpenStudy (janu16):

@nincompoop can you please help?

OpenStudy (janu16):

@perl

geerky42 (geerky42):

by plug in x=3

geerky42 (geerky42):

Just replace x to 3.

OpenStudy (janu16):

so it would be x=4(y+3)-5 x=4(3+3)-5?

OpenStudy (janu16):

@geerky42

OpenStudy (danjs):

Are you still working on this ?

OpenStudy (janu16):

yes

OpenStudy (danjs):

f(x) = y = 4(x+3) - 5 is a line that is given y = 4x + 12 -5 y = 4x + 7

OpenStudy (janu16):

ya so when do you substitute x for 3?

OpenStudy (danjs):

The inverse of that , will be another line that is reflected over the line with slope 1, y=x. To figure that , you can just switch x and y around and resolve for the new y.

OpenStudy (danjs):

Original: y = 4x + 7 Inverse: x = 4y + 7 x - 7 = 4y y = (x-7) / 4

OpenStudy (danjs):

Put x=3 in the inverse function, calculate the y value

OpenStudy (janu16):

-1?

geerky42 (geerky42):

That's correct.

OpenStudy (danjs):

See it is a reflection if you switch the x and y around, as in the graph

OpenStudy (janu16):

ya i see it. it really helped thank you so much and do you know how to do f(g(x) problems?

OpenStudy (danjs):

sure

OpenStudy (janu16):

well i need help with Given that f(x) = x2 + 2x + 3 and g(x) = quantity of x plus four, over three, solve for f(g(x)) when x = 2.

OpenStudy (janu16):

like i dont know how to setup

OpenStudy (danjs):

solve g(2) first, what is that

OpenStudy (danjs):

g(x) = (x+4) / 3

OpenStudy (janu16):

2?

OpenStudy (janu16):

wait do you put 2 for x?

OpenStudy (danjs):

yes, now find f[g(2)], what do ya think to do?

OpenStudy (janu16):

4x2x+3 than 6x=3 which is 2?

OpenStudy (danjs):

since g(2) = 2 , start from the middle and work outwards f(2) = 2^2 + 2*2 + 3

OpenStudy (janu16):

11?

OpenStudy (danjs):

yeah

OpenStudy (janu16):

oh i forgot to substitute x with 2 for 2x

OpenStudy (janu16):

so if i haveanother problem like in future do i solve g first? or depending on the problem

OpenStudy (danjs):

for compound functions like f([g(x)] start with the inner function, evaluate if given a value, then use g(x) value in f(x)... work inner function to outer functions

OpenStudy (janu16):

ohok thank you very much!

OpenStudy (danjs):

welcome, goodluck

OpenStudy (janu16):

there are like some problems like f(x) and g(x) and you have to compare them. there are lot of problems but i dont know how to compare like if you know how to do it i will put it on new page

OpenStudy (danjs):

sure, just tag my name with @name

OpenStudy (janu16):

ok thanks